OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a19 = -58, a21 = -164

OpenStudy (anonymous):

@amistre64 do you mind helping me with this one too? would I use the formula an=dn+c?

OpenStudy (amistre64):

something simliar yes

OpenStudy (amistre64):

\[a_n=a_1+d(n-1)\]

OpenStudy (amistre64):

let a1 = a19, then a3 = a21

OpenStudy (amistre64):

-164=-58 +d(3-1); solve for d

OpenStudy (anonymous):

so -53?

OpenStudy (amistre64):

yes therefore: a19 = a1 -53(18); solve for a1

OpenStudy (anonymous):

ahh okay 896

OpenStudy (amistre64):

therefore: an = 896 -53(n-1)

OpenStudy (anonymous):

Okay awesome! thank you for the right formula!

OpenStudy (amistre64):

yep